# Understanding Radiation Flow in a Stochastic Medium

### v1

@@ 4.3 Multi-Clump Studies: Putting it all Together
Realistic inhomogeneous radiation-flow problems typically involve chaotic inhomogeneities with many dense clumps. To understand better how radiation flows through a multi-clump medium, we expand our study to include a series of calculations with 25 clumps in our flow region (Byvank et al., 2024) varied the size and number of clumps in a similar experimental setup). To better understand the additional complexity caused by this structure, we have run a series of randomly spaced multi-clump calculations. We continue to use the clump properties (density, composition) from the XFOL and XFLOWS laboratory experiments (Johns et al., 2023; Byvank et al., 2024). We then vary the number, size and distribution of the clumps to better understand the broader implications of inhomogeneous radiation flow. In particular, these studies allow us to investigate the effects of blow-off (and its dependence on the radiation pressure) under more realistic conditions.
For our first study, we focus on the impact of hydrodynamic feedback on the flow of radiation. Figure 12 shows the temperature profile of the same initial clump profile, but with two different drives, each comparing a simulation that is pure transport (“no hydro”)) with another that includes hydrodynamic feedback. The pure-transport calculations show an extreme limit where the radiation simply flows around the clumps. For the lower drives, the hydrodynamic feedback is faster than the flow timescale and the clumps expand to block the radiation flow. The radiation flow is dramatically altered by the blow-off from the clumps. With a stronger drive, the outflow is minimized due to both higher radiation pressures and the fact that the flow is faster, allowing less time for hydrodynamic feedback. In the stronger-drive simulations, the differences between the pure-transport and radiation-hydrodynamics simulations are less dramatic.
Figure 12: Temperature maps for four cylindrical models with 25 material clumps randomly spaced in the flow region. We use our standard clump and flow densities and compare the solution with and without hydrodynamics turned on. We also vary the temperature of the drive. At our standard temperatures, radiative feedback dramatically alters the flow, but, as we increase the temperature, the differences between pure radiation (“no hydro”) and radiation-hydrodynamics simulations become smaller.
@@
Figure 13: Distribution of radiation-flow timescales for 100 simulations of 25-clump simulations using our standard clump properties ($\rho=6.1{\rm\,g\,cm^{-3},10\,\mu m\,diameter}$), drive power and flow region conditions. The two distributions correspond to the timescale required for a temperature of 25% or 75% of the radiation front at 0.05 cm above the clumpy medium to rise above 20 eV. The spread in the distribution is high (over 2 ns, $>$ 20%) and, at least for these 100 simulations, does not appear to be well-matched by a Gaussian.
@@
Figure 14: Material and radiation temperature for two 100-clump calculations where the radiation temperature drive is 1 keV. The two calculations use two different clump sizes ($r_{\rm clump}=0.05\mu m,0.1\mu m$). The corresponding clump covering fractions of these two models are 5 and 20%. We plot the radiation fronts as they reach the edge of our simulation grid, at 7.8 ps and 28 ps for the small and large clumps, respectively.
@@
Figure 16: Temperature of a strong-drive radiation flow through a clumpy medium where the clump radius is 15 $\mu$m and the number of clumps varies from 25 to 98. As the covering fraction increases, the radiation timescale for the propagation of the radiation front increases dramatically and the times for each panel are very different.
@@
Figure 17: Temperature of a strong-drive radiation flow through a clumpy medium where the clump radius and number of clumps are modified to produce the same covering fraction. Comparing to Figure 16 demonstrates the importance of the clump size on the radiation flow.
Clearly, even in the radiation-dominated case, hydrodynamics effects like outflow/winds will be important. The subgrid formulae in Section 3.1 will struggle to match this data. The Sundqvist et al. (2018) formulae discussed in Section 3.1 do not reproduce the magnitude of this variation. Although the additional scattering term in the Olson (2007) approach can come closer to fitting the data, this requires incorporating the temperature-dependence of the opacity. We defer a detailed application-specific comparison of these subgrid models to a later paper.
5 Applications Revisited
@@
Radiation flow through inhomogeneous media is critical for a broad range of problems in astrophysics and beyond (see Table 1). In many cases, the inhomogeneities can not be resolved in the large-scale calculations of these applications. In this paper, we reviewed some of the recipes developed to capture radiation transport through these media and the experiments developed to test them.
@@
This paper included a large set of simulations to better understand the physics behind radiation flow through an inhomogeneous medium. We found that for regimes where the radiation pressure is not dominant, radiative heating of the high-density clumps can cause the clumps to expand. This wind can increase the opacity throughout the medium, dramatically altering the flow of radiation. Most of the recipes in the literature focus on transport-only effects, which are more valid for the conditions where radiation pressure dominates. But, as we have seen in this paper, even in a radiation-pressure dominated system, radiation-hydrodynamics effects can not be ignored.
@@
In all of our models, energy deposition from the radiation front played a much bigger role than momentum deposition and hence, “outflows” or “winds” were more important for the subsequent evolution than ablation. For less dense clumps, as we might expect from turbulent instabilities, ablation can play a much more important role Fryer et al. (2020a). Ablation will have a similar effect to blow-off in the sense that it will disperse the clumpy material, causing shocks and cutting off the flow. But the nature of ablation will be different than the models shown here and much more work must be done to understand all aspects of inhomogeneous radiation flow for the broad set of applications discussed in this paper.
@@
The work by CLF was supported by the US Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). A portion of the work by CLF was performed at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611.

### v3

@@ 4.3 Multi-Clump Studies: Putting it all Together
Realistic inhomogeneous radiation-flow problems typically involve chaotic inhomogeneities with many dense clumps. To understand better how radiation flows through a multi-clump medium, we expand our study to include a series of calculations with 25 clumps in our flow region (Byvank et al., 2024) varied the size and number of clumps in a similar experimental setup). The simulations shown here all have a size of 0.04 cm in x-dimension and a y-dimension size of 0.08 cm. For our base model, clumps have radii of $50\mu m$. The mean free path in the flow region ($\lambda_{\rm mfp}^{\rm flow}$) increases as the material heats but is roughly 0.005-0.01 cm and the mean free path in the clumps is ($\lambda_{\rm mfp}^{\rm clump}$) is $1-2\mu m$. We also vary the clump size to study the importance of the flow mean free path to the clump size. To better understand the additional complexity caused by this structure, we have run a series of randomly spaced multi-clump calculations. We continue to use the clump properties (density, composition) from the XFOL and XFLOWS laboratory experiments (Johns et al., 2023; Byvank et al., 2024). We then vary the number, size and distribution of the clumps to better understand the broader implications of inhomogeneous radiation flow. In particular, these studies allow us to investigate the effects of blow-off (and its dependence on the radiation pressure) under more realistic conditions.
For our first study, we focus on the impact of hydrodynamic feedback on the flow of radiation. Figure 12 shows the temperature profile of the same initial clump profile, but with two different drives (our standard drive and one where we multiply the temperature of the radiation-driven drive by a factor of 2), each comparing a simulation that is pure transport (“no hydro”)) with another that includes hydrodynamic feedback. The pure-transport calculations show an extreme limit where the radiation simply flows around the clumps. For the lower-temperature drives, the hydrodynamic feedback is faster than the flow timescale and the clumps expand to block the radiation flow. The radiation flow is dramatically altered by the blow-off from the clumps. With a stronger drive, the outflow is minimized due to both higher radiation pressures and the fact that the flow is faster, allowing less time for hydrodynamic feedback. In the stronger-drive simulations, the differences between the pure-transport and radiation-hydrodynamics simulations are less dramatic.
Figure 12: Effective radiation temperature maps for four cylindrical models with 25 material clumps randomly spaced in the flow region. We use our standard clump and flow densities and compare the solution with and without hydrodynamics turned on. We also vary the temperature of the drive. At our standard temperatures, radiative feedback dramatically alters the flow, but, as we increase the temperature, the differences between pure radiation (“no hydro”) and radiation-hydrodynamics simulations become smaller. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
@@
Figure 13: Distribution of radiation-flow timescales for 100 simulations of 25-clump simulations using our standard clump properties ($\rho=6.1{\rm\,g\,cm^{-3},10\,\mu m\,diameter}$), drive power and flow region conditions. The two distributions correspond to the timescale required for a temperature of 25% or 75% of the radiation front at 0.05 cm above the clumpy medium to rise above 20 eV. The spread in the distribution is high (over 2 ns, >> 20%) and, at least for these 100 simulations, does not appear to be well-matched by a Gaussian.
@@
Figure 14: Material and effective radiation temperature for two 100-clump calculations where the radiation temperature drive is 1 keV. The two calculations use two different clump sizes ($r_{\rm clump}=0.05\mu m,0.1\mu m$). The corresponding clump covering fractions of these two models are 5 and 20%. We plot the radiation fronts as they reach the edge of our simulation grid, at 7.8 ps and 28 ps for the small and large clumps, respectively. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
@@
[…]
@@
Figure 17: Effective radiation temperature of a strong-drive radiation flow through a clumpy medium where the clump radius and number of clumps are modified to produce the same covering fraction. Comparing to Figure 16 demonstrates the importance of the clump size on the radiation flow. Our code models the radiation with 71 energy groups. The effective radiation temperature is inferred by fitting the spectrum from these groups.
Clearly, even in the radiation-dominated case, hydrodynamics effects like outflow/winds will be important. The subgrid formulae in Section 3.1 will struggle to match this data. To test these models, we created a simulation with a simplified opacity prescription, $\sigma_{\rm flow}=10^{3}$, $\sigma_{\rm clump}=10^{4}$, within our 100 clump calculation. The density of the clumps is $6.1\,{\rm g\,cm^{-3}}$ and the flow region is $0.1\,{\rm g\,cm^{-3}}$. The corresponding effective opacity for the “Olson” (Prinja & Olson, 2005; Olson, 2007) formulation yielded $\sigma_{a,eff}=1046.2$ (equation 1) and $\sigma_{s,eff}=2561$ (equation 2. The corresponding opacity from the “Owocki and Sundqvist” (Owocki & Sundqvist, 2018) prescription is $\sigma_{eff}=3949$ (equation 3). Figure 18 shows radiation flow calculations at 2 ns for the clumpy run and these two prescriptions. For this calculation, a simple mass-mixed model (uniformly mixing, by mass, the clump material in the transport region) is a better fit to the data.
Figure 18: Radiation temperature at 2 ns for a 100 clump simulation using constant opacities: $\sigma_{\rm flow}=10^{3}$, $\sigma_{\rm clump}=10^{4}$ with corresponding densities of $\rho_{\rm clump}=6.1\,{\rm g\,cm^{-3}}$and $\rho_{\rm flow}=6.1\,{\rm g\,cm^{-3}}$. This simulation is compared to the radiation propagation for both the “Olson” (equations 12) formulation and the “Owocki and Sundqvist” prescription (equation 3). A mass-mixed model (uniformly mixing the clump material throughout the transport regime) is also included in the figure. This latter solution is a much closer fit to the full solution, likely because the hydrodynamic reaction to the radiation flow causes considerable mixing.
It is not surprising that the formulations discussed in Section 3.1 are not ideal fits to a radiation hydrodynamics solution. Prinja & Olson (2005); Olson (2007) clearly state that they are testing their prescription to a pure transport model. And, as we have shown, hydrodynamics effects can be extremely important. If we turn hydrodynamics off, we find in Figure 19 that our prescriptions are a much better fit to the full calculations. A much more extensive study is required to truly understand the full physics of radiation flow through a multi-clump medium. We defer this more extensive study, along with a detailed application-specific comparison of these subgrid models to a later paper.
Figure 19: Radiation temperature at 2 ns for the same simulation as Figure 18. In this case, the clumpy medium is modeled with the hydrodynamics turned off. Both of our prescriptions, “Olson” and “Owocki and Sundqvist” are much better fits to this solution.
5 Conclusions
@@
Radiation flow through inhomogeneous media is critical for a broad range of problems in astrophysics and beyond (see Table 1). In many cases, the inhomogeneities can not be resolved in the large-scale calculations of these applications. In this paper, we reviewed some of the recipes developed to capture radiation transport through these media and the experiments developed to test them. Existing subgrid models designed to capture unresolved clumps mostly focus on altering the opacity based on the filling factor and opacity of the clumps. We find that these prescriptions do not match the detailed simulations when hydrodynamic feedback is included. As we have shown in this paper, the base opacity is also affected by the specific heat of the clumps (because it alters the blow-off) and the size-scale of the clumps. Any recipe ultimately should include the stochasticity in the final result as well (recall the variation in Figure 13).
@@
This paper included a large set of simulations to better understand the physics behind radiation flow through an inhomogeneous medium. We found that for regimes where the radiation pressure is not dominant, radiative heating of the high-density clumps can cause the clumps to expand. This wind can increase the opacity throughout the medium, dramatically altering the flow of radiation. Most of the recipes in the literature focus on transport-only effects, which are more valid for the conditions where radiation pressure dominates. But, as we have seen in this paper, even in a radiation-pressure dominated system, radiation-hydrodynamics effects can not be ignored. In many supernova-interaction problems, this blow-off will play an important role in defining both the broad energy profile and line features of the supernova spectra (Fryer et al., 2020a; Rastinejad et al., 2025). Radiation will drive interactions prior to the interaction of material shocks and the blow-off material can be more easily swept up in the supernova blastwave. Several supernova groups are now developing radiation-hydrodynamics codes with higher-order radiation transport schemes to study these effects.
@@
In all of our models, energy deposition from the radiation front played a much bigger role than momentum deposition and hence, “outflows” or “winds” were more important for the subsequent evolution than ablation. For less dense clumps, as we might expect from turbulent instabilities, ablation can play a much more important role (Fryer et al., 2020a). Ablation will have a similar effect to blow-off in the sense that it will disperse the clumpy material, causing shocks and cutting off the flow. Ablation may be more important in stellar winds and protostellar outflows. But the nature of ablation will be different than the simulations shown here and much more work must be done to understand all aspects of inhomogeneous radiation flow for the broad set of applications discussed in this paper.
In most cases, the hydrodynamic feedback effects are not fully captured by current prescriptions designed to capture a modified total opacity or one that incorporates an effective scattering term to capture angular effects. Especially in problems studying detailed spectra (e.g. in laboratory experiments or supernova light-curves), the blow-off mixing and shock interactions must also be captured by any subgrid prescription.
The experiments can be developed to produce scalable rations of the mean free path ($\lambda_{\rm mfp}^{\rm flow}$), clump mean free path ($\lambda_{\rm mfp}^{\rm clump}$) and clump size ($d^{\rm clump}$): e.g. $\lambda_{\rm mfp}^{\rm flow}/\lambda_{\rm mfp}^{\rm clump}$, $\lambda_{\rm mfp}^{\rm flow}\rm/d^{\rm clump}$. These allow us to probe the radiation flow in transport conditions that mimic those of a broad range of applications. In addition, these experiments can study our numerical models to couple radiation and hydrodynamics to determine the uncertainties due to numerical effects like numerical diffusion (both in matter and radiation).
@@
This work was supported by the US Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). A portion of the work by CLF was performed at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611.